OpenStudy (anonymous):

Find f'(x) of f(x)=1/x^2

hartnn (hartnn):

know the derivative of x^n ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

then 1/x^2 is just x^-2 plug in n=-2 in that formula

OpenStudy (anonymous):

Yeah I know that but I'm suppose to use the formula f'(x)=h-> 0f(x+h)-f(x)/h

hartnn (hartnn):

ok, then whats f(x+h) = .. ?

OpenStudy (anonymous):

1/(x+h)^2

hartnn (hartnn):

yes, so, your numerator f(x+h) - f(x) will be 1/ (x+h)^2 + 1/x^2 try to simplify this ? by making a common denominator?

hartnn (hartnn):

soory, there should be a minus - in between

hartnn (hartnn):

1/ (x+h)^2 - 1/x^2

OpenStudy (anonymous):

umm would it be x^2/(x+h)^2x^2 - (x+h)^2/x^6(x+h)^2/h

hartnn (hartnn):

how is it x^6 there ? did you mean x^2 ? everything else seem to be correct

hartnn (hartnn):

so we have a common denominator of x^2 (x+h)^2 now , right ? when we combine the numerator, we get \(\Large \dfrac{x^2-(x+h)^2}{h x^2(x+h)^2}\) did you get this ?

OpenStudy (anonymous):

oh sorry x^2

OpenStudy (anonymous):

yes i did

hartnn (hartnn):

simplify the numerator, anything getting cancelled ?

OpenStudy (anonymous):

umm the x^2?

hartnn (hartnn):

correct! from whatever remains, factor out the "h"

OpenStudy (anonymous):

so h(2x+h)/hx^2(x+h)^2?

hartnn (hartnn):

there will be a minus sign, right ? \(\Large \dfrac{x^2-x^2 -2xh-h^2}{h x^2(x+h)^2}= \Large \dfrac{-\cancel h (2x+h)}{\cancel h x^2(x+h)^2}\) got this ?

hartnn (hartnn):

got how that "h" got cancelled ?

OpenStudy (anonymous):

ohh yesss i see lol i forgot the minus sign

OpenStudy (anonymous):

so what happens to the h in the denominator?

hartnn (hartnn):

one of the 'h' got cancelled, right now since we have limit as h->0 we can directly plug in h =0 in the function now! :) what do u get ?

OpenStudy (anonymous):

2x?

hartnn (hartnn):

numerator = -2x yes, what about the denominator, when h= 0 ?

OpenStudy (anonymous):

hmm im not sure..

hartnn (hartnn):

the denominator is x^2 (x+h)^2 right ? so what u get when u put h= 0 ?

OpenStudy (anonymous):

x^4?

hartnn (hartnn):

yes! so we have -2x/x^4 what does that simplify to ?

OpenStudy (anonymous):

-2x^-3?

hartnn (hartnn):

absolutely correct! derivative of 1/x^2 is indeed -2/x^3 or -2x^-3

OpenStudy (anonymous):

yay! thank you so so much you made this 10 times easier to understand ^-^

hartnn (hartnn):

i am glad to hear that :) you're most welcome ^_^